OpenStudy (anonymous):

Part A: DO NOT SOLVE! You are only setting up a linear system! (40 points) The table below shows the number of calories burned for a 130-lb person in a minute. If you want to work-out for a total of 60 minutes and burn 500 calories, choose two activities and set up a linear system. Activity Calories Burned per Minute Bicycling 10 Hiking 7 Running 11 Walking 5 Swimming 12 • Let x = total minutes for activity 1 and y = total minutes for activity 2. • Equation 1 should represent total minutes. Equation 2 should represent total calories burned. Equation 1: Equation 2: help

OpenStudy (anonymous):

Wait where is the table ?

OpenStudy (anonymous):

the activities are the table

OpenStudy (anonymous):



OpenStudy (anonymous):

Which are you using? which activity?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Which activity are you planning on using for the equations?

OpenStudy (anonymous):

i can use any of them

OpenStudy (anonymous):

biking and swimming.

OpenStudy (amistre64):

then calories for B plus calories for S

OpenStudy (amistre64):

b+ s = 60 10b + 12s = 500

OpenStudy (anonymous):

10 and 12

OpenStudy (anonymous):

22

OpenStudy (anonymous):

ill be quiet.

OpenStudy (amistre64):

minutes biking, plus minutes swimming equals 60 minutes total 10calories per bike minute plus 12 calories per swim minute = 500 calories

OpenStudy (amistre64):

solve the system of equations: b+ s = 60 10b + 12s = 500

OpenStudy (anonymous):

We just have to leave it in a equation

OpenStudy (amistre64):

lol, yeah i just noticed that :)

OpenStudy (amistre64):

sigh, i gave the final answer without realisng it

OpenStudy (anonymous):

sorry but thank you for everything (: @amistre64

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

don't sigh. its okay